OpenStudy (anonymous):

in the last two years, the average math score at a local high school increased to 90 from 72. by what percent did the scores increase? a) 2.5% b) 18% c) 25% d) 162%

OpenStudy (carson889):

90-72 = 18 18/72 = 9/36 = 3/12 = 1/4 = 0.25 * 100% = 25% Verify: 72 * 1.25 = 90 72 + 72*.25 = 72 + 18 = 90

OpenStudy (anonymous):

what you do i you subtraction and then divide bu two